Ʌ0 Al2(SO4)3 is 858 scm2 mole-1 . Find Ʌ0 of Al3+ if Ʌ0 of SO4 2- is 160 scm2 mole-1 . 

Grade:12

1 Answers

Vikas TU
7 years ago
K = ks1*ks2 = given for Al2(SO4)3
Hence,
858 Scm2/mol = 160*Ks2
ks2 = 5.3625 would be the  Ʌ0 for Al3+